OpenStudy (anonymous):

Given the function g(x) = −4x + 5, find g(−4). −11 21 2.25 11

OpenStudy (anonymous):

g(-4) = substitute all x's in -4x+5 with (-4) and simplify

OpenStudy (anonymous):

so it would be g(-4) = -4(-4) +5 = 16 +5 = 21